OpenStudy (anonymous):

A line passes through the origin and has a slope of 4. What is the equation of the line that is parallel to the first line and passes through the point (1,-2)? A. y=4x+9 B. y=-4x+2 C. 4x-6 D. y=-1/4x-7/4 A line passes through the origin and has a slope of 4. What is the equation of the line that is parallel to the first line and passes through the point (1,-2)? A. y=4x+9 B. y=-4x+2 C. 4x-6 D. y=-1/4x-7/4 @Mathematics

OpenStudy (anonymous):

To get this we'll just use point-slope form of the line: y-y1 = m(x-x1) From the information given, m = 4, y1 = -2, x1 = 1 y + 2 = 4(x-1) y+2 = 4x-4 y = 4x-6